An electric lamp of mass 900g hangs on a flex, which is pulled aside by a horizontal string until the flex makes an angel of 30° by with the vertical. Find the tension in the string and in the flex?
1
Expert's answer
2020-01-27T09:30:48-0500

The tension in the flex:


"T_1=\frac{mg}{\cos{30}}=\frac{(0.9)(9.8)}{\cos{30}}=10.2\ N"

The tension in the string



"T_2=mg\tan{30}=(0.9)(9.8)\tan{30}=5.1\ N"